OpenStudy (anonymous):

Consider the following equations:

OpenStudy (shubhamsrg):

I will consider them if I could see them :/

OpenStudy (anonymous):

1) 1=1 2) 2+3+4=1+8 3) 5+6+7+8+9=8+27